# Removing Undesirable Concepts in Text-to-Image Generative Models with Learnable Prompts

### v1

@@ Abstract
Generative models have demonstrated remarkable potential in generating visually impressive content from textual descriptions. However, training these models on unfiltered internet data poses the risk of learning and subsequently propagating undesirable concepts, such as copyrighted or unethical content. In this paper, we propose a novel method to remove undesirable concepts from text-to-image generative models by incorporating a learnable prompt into the cross-attention module. This learnable prompt acts as additional memory to transfer the knowledge of undesirable concepts into it and reduce the dependency of these concepts on the model parameters and corresponding textual inputs. Because of this knowledge transfer into the prompt, erasing these undesirable concepts is more stable and has minimal negative impact on other concepts. We demonstrate the effectiveness of our method on the Stable Diffusion model, showcasing its superiority over state-of-the-art erasure methods in terms of removing undesirable content while preserving other unrelated elements.
**footnotetext: Equal contribution where Anh Bui is the corresponding author.
@@ 1 Introduction
Recent advances in text-to-image generative models (Rombach et al., 2022; Ramesh et al., 2021, 2022) have captured significant attention due to their outstanding image quality and boundless creative potential. These models undergo training on extensive internet datasets, equipping them with the ability to replicate a diverse array of concepts. Nevertheless, because of the abundance of undesirable concepts in the training data, such as racism, sexism, and violence, these models can learn and propagate these concepts, which can be exploited by users to generate harmful content, contributing to the proliferation of fake news, hate speech, and disinformation (Rando et al., 2022; Qu et al., 2023; Westerlund, 2019). Therefore, excluding these undesirable concepts from the model’s output is a critical step in ensuring the safety and usefulness of these models.
Numerous prior strategies have been explored to address this challenge, including dataset filtering (StabilityAI, 2022), post-generation filtering (Rando et al., 2022), and inference guiding (Schramowski et al., 2023). Dataset filtering entails the heavy and time-consuming task of removing objectionable data from the training dataset, which proves especially taxing for large models and datasets. Additionally, this process has been observed to lead to a decline in output quality (StabilityAI, 2022). Post-generation filtering, which involves implementing classifiers like the NSFW (Not-Safe-For-Work) classifier (StabilityAI, 2022) to sift out sensitive content after generation, is hindered by its limited capacity to effectively detect such content while maintaining a low false-positive rate (Rando et al., 2022). Additionally, users can easily disable the NSFW classifier, making it an unreliable safeguard. On the contrary, the introduction of guidance during the inference process, as demonstrated in Schramowski et al. (2023), offers a simple yet effective means to steer clear of generating sensitive content without the need for model retraining. However, it has also been exposed to possible bypass by users with access to the model parameters (Gandikota et al., 2023). It is desirable to truly erase undesirable concepts from the models themselves. This necessitates model retraining with appropriate loss functions and mechanisms designed to eliminate predefined undesirable concepts while still preserving and maintaining others.
Some approaches have been proposed to retrain foundation models to eliminate undesirable concepts, typically Gandikota et al. (2023); Kumari et al. (2023); Orgad et al. (2023); Zhang et al. (2023). Specifically, these approaches aim to directly modify the parameters of the foundation models to optimize appropriate losses. However, the parameter space is shared among all concepts, and semantic-related concepts often trigger highly overlapping sets of parameters. Therefore, erasing an undesirable concept might significantly affect its semantically relevant retaining concepts. To address this issue, inspired by the success of parameter prompt-tuning approaches (Li & Liang, 2021; Lester et al., 2021; Pfeiffer et al., 2020), we incorporate a parameter prompt serving as additional memory to the cross-attention layers of foundation models. These additional parameter prompts are first trained to effectively generate undesirable concepts, aiming to reduce the dependency of undesirable concepts on the parameters of foundation models and corresponding textual inputs. Additionally, this helps minimize modifications to the parameters of foundation models when eliminating undesirable concepts later, as knowledge of undesirable concepts is transferred to the additional memory induced by the parameter prompts. In the subsequent step, with the aid of the additional prompt, we actively fine-tune the parameters of foundation models to erase undesirable concepts by enforcing the output associated with these concepts close to a neutral one. The experiments, conducted in three settings— ➀ erasing object-oriented concepts, ➁ mitigating unethical content, and ➂ erasing artistic style concepts—show that our proposed approach outperforms the state-of-the-art erasure methods in both erasing undesirable concepts and preserving the retaining concepts.
@@ 2 Related Work
In this section, we provide an extensive overview of existing literature concerning concept erasure and related techniques pertinent to our work.
@@
Pre-processing methods represent a straightforward approach to eliminating undesired concepts from input images. This involves employing pre-trained detectors to identify images containing objectionable content and subsequently excluding them from the training set. However, the drawback lies in the necessity of retraining the model from scratch, which proves computationally expensive and impractical for evolving erasure requests. A notable instance of complete retraining is evident in Stable Diffusion v2.0 (StabilityAI, 2022), but this approach was reported to leave the model inadequately sanitized (Gandikota et al., 2023).
Post-processing methods encompass the utilization of Not-Safe-For-Work (NSFW) detectors to identify potentially inappropriate content in generated images. Images flagged by the NSFW detector are then either blurred or blacked out before being presented to users. This method, employed by organizations such as OpenAI (developer of Dall-E), StabilityAI (developer of Stable Diffusion), and Midjourney Inc (developer of Midjourney), is considered highly effective. However, the open-source nature of the Stable Diffusion model exposes it to potential evasion by modifying the NSFW detector in the source code. Closed-source models, like Dall-E, are not immune either, as demonstrated in (Yang et al., 2024), where a technique similar to Boundary Attack (Brendel et al., 2017) was used to uncover adversarial prompts that could bypass the filtering mechanism.
Concept Mimicry serves as a personalization technique, generating images aligned with a user’s preferences based on their input. Noteworthy methods include Textual Inversion (Gal et al., 2022) and Dreambooth (Ruiz et al., 2023), which have proven effective with minimal user input. In contrast, Anti Concept Mimicry is employed to safeguard personal or artistic styles from being copied through Concept Mimicry. Achieved by introducing imperceptible adversarial noise to input images, this technique can deceive Concept Mimicry methods under specific conditions. Recent contributions such as Anti-Dreambooth (Van Le et al., 2023) have explored and demonstrated the effectiveness of this approach.
To date, the most successful strategy for sanitizing open-source models, such as Stable Diffusion, involves cleaning the generator (e.g., U-Net) in the diffusion model post-training on raw, unfiltered data and before public release. This approach, as partially demonstrated in Gandikota et al. (2023), underscores the importance of addressing potential biases and undesired content in models before their deployment.
@@
Existing erasing methods (Gandikota et al., 2023; Orgad et al., 2023; Zhang et al., 2023; Kumari et al., 2023) aim to erase undesirable concepts by fine-tuning foundation models with appropriate losses to unlearn and erase these undesirable concepts. Specifically, TIME (Orgad et al., 2023), UCE (Zhang et al., 2023), Concept Ablation (Kumari et al., 2023), and SDD try to project meaning of harmful context into another benign one, while ESD (Gandikota et al., 2023) uses the principle of classifier-free-guidance to remove the distribution of the bad concept from the LDM.
@@
The overarching concept behind prompting involves applying a function to alter the input text, providing the language model with supplementary task-related information. However, devising an effective prompting function presents challenges and necessitates heuristic approaches. Recent studies, such as prompt tuning (Lester et al., 2021) and prefix tuning (Li & Liang, 2021), attempt to tackle this challenge by employing trainable prompts within a continuous space, resulting in impressive performance in transfer learning tasks. Prompts encapsulate task-specific knowledge with significantly fewer additional parameters compared to competing methods like Adapter (Pfeiffer et al., 2020; Wang et al., 2021) and LoRA (Hu et al., 2022).
3 Background
3.1 Diffusion based Text-to-Image Generative Models
Denoising Diffusion Models:
Generative modeling is a fundamental task in machine learning that aims to approximate the true data distribution $p_{\text{data}}$ from a dataset $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{N}$. Diffusion models, a recent class of generative models, have shown impressive results in generating high-resolution images (Ho et al., 2020; Rombach et al., 2022; Ramesh et al., 2021, 2022). In a nutshell, training a diffusion model involves two processes: a forward diffusion process where noise is gradually added to the input image, and a reverse denoising diffusion process where the model tries to predict a noise $\epsilon_{t}$ which is added in the forward process. More specifically, given a chain of $T$ diffusion steps $x_{0},x_{1},...,x_{T}$, the denoising process can be formulated as follows:
$p_{\theta}(x_{T:0})=p(x_{T})\prod_{t=T}^{1}p_{\theta}(x_{t-1}\mid x_{t})$ (1)
The model is trained by minimizing the difference between the predicted noise $\epsilon_{t}$ and the true noise $\epsilon$ as follows:
$\mathcal{L}=\mathbb{E}_{x_{0}\sim p_{\text{data}},t,\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left\|\epsilon-\epsilon_{\theta}(x_{t},t)\right\|_{2}^{2}$ (2)
where $\epsilon_{\theta}(x_{t},t)$ is the predicted noise at step $t$ by the denoising model $\theta$.
Latent Diffusion Models:
With an intuition that semantic information that controls the main concept of an image can be represented in a low-dimensional space, (Rombach et al., 2022) proposed a diffusion process operating on the latent space to learn the distribution of the semantic information which can be formulated as follows:
$p_{\theta}(z_{T:0})=p(z_{T})\prod_{t=T}^{1}p_{\theta}(z_{t-1}\mid z_{t})$ (3)
where $z_{0}\sim\varepsilon(x_{0})$ is the latent vector obtained by a pre-trained encoder $\varepsilon$.
The objective function of the latent diffusion model as follows:
$\mathcal{L}=\mathbb{E}_{z_{0}\sim\varepsilon(x),x\sim p_{\text{data}},t,\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left\|\epsilon-\epsilon_{\theta}(z_{t},t)\right\|_{2}^{2}$ (4)
3.2 Conditioning Cross-Attention
Working on the latent space allows the diffusion process to be conditioned on a textual description more naturally than working on the image space. It can be achieved by using a cross-attention module (Vaswani et al., 2017) which allows the model to attend and align the textual content from the input description to the visual feature from the latent representation $\mathbf{z}$. More specifically, at a specific cross-attention layer i of the U-Net architecture, the query input matrix is a batch of $b$ latent representations from the previous diffusion step denoted as $\mathbf{Z}\in\mathbb{R}^{[b,m_{z},d_{z}]}$, where $m_{z}$ is the spatial dimension of the latent representation $z$ (i.e., $m_{z}=h\times w$) and $d_{z}$ is the dimension of the vector representation of each spatial location. The key and value input matrices are the same, which is a batch of $b$ corresponding textual embeddings denoted as $\mathbf{C}\in\mathbb{R}^{[b,m_{c},d_{c}]}$, where $m_{c}$ is the sequence length and $d_{c}$ is the size of the token embedding vector.
The cross-attention module outputs $\mathbf{O}\in\mathbb{R}^{[b,m_{z},d]}$ by following the Query-Key-Value (QKV) mechanism (Vaswani et al., 2017) as summarized in Table 1.
Table 1: Cross-Attention Mechanisms. $\text{cat}(),\text{repeat}(\cdot,b),\sigma()$ represent the concatenate (at dim=1), repeat an input $b$ times (at dim=0) and the softmax operations (at dim=2), respectively.
Original Concatenative Addititive
Operation Dim Operation Dim Operation Dim
$Q$ $W_{q}Z$ $b\times m_{z}\times d$ $W_{q}Z$ $b\times m_{z}\times d$ $W_{q}Z$ $b\times m_{z}\times d$
$K$ $W_{k}C$ $b\times m_{c}\times d$ $W_{k}\;\text{cat}(C,\text{repeat}(p,b))$ $b\times(m_{c}+m_{p})\times d$ $W_{k}\;\text{cat}(C+\text{repeat}(p,b))$ $b\times m_{c}\times d$
$V$ $W_{v}C$ $b\times m_{c}\times d$ $W_{v}\;\text{cat}(C,\text{repeat}(p,b))$ $b\times(m_{c}+m_{p})\times d$ $W_{v}\;\text{cat}(C+\text{repeat}(p,b))$ $b\times m_{c}\times d$
$A$ $\sigma(QK^{T}/\sqrt{d})$ $b\times m_{z}\times m_{c}$ $\sigma(QK^{T}/\sqrt{d})$ $b\times m_{z}\times(m_{c}+m_{p})$ $\sigma(QK^{T}/\sqrt{d})$ $b\times m_{z}\times m_{c}$
$O$ $AV$ $b\times m_{z}\times d$ $AV$ $b\times m_{z}\times d$ $AV$ $b\times m_{z}\times d$
@@ 4.1 Motivation
Erasing a concept from a foundation model is not a difficult task. As demonstrated in prior works (Gandikota et al., 2023; Kumari et al., 2023; Orgad et al., 2023; Zhang et al., 2023), fine-tuning the foundation model to enforce the output associated with the to-be-erased concept to be close to the output with a neutral or null concept, can effectively remove the undesirable concept. However, the parameter space is shared among all concepts, and semantic-related concepts often trigger highly overlapping sets of parameters. Therefore, the erasure effect commonly comes with a significant drop in the model’s capability in other concepts. As later revealed in Section 5.5, erasing an undesirable concept, even with state-of-the-art methods, still affects the model’s semantically relevant retaining concepts. This is evident in the instability of the alignment between visual and textual features in neutral or retained concepts along the fine-tuning process.
To tackle this challenge, drawing inspiration from the success of parameter prompt-tuning approaches in Natural Language Processing (Li & Liang, 2021; Lester et al., 2021; Pfeiffer et al., 2020), we introduce a learnable parameter prompt as additional memory to the cross-attention layers of the foundation model. Intuitively, our method involves two alternate processes named knowledge transfer and knowledge removal, with the additional prompt acting as a buffer between them. In the knowledge transfer stage, the prompt is trained to effectively generate undesirable concepts, aiming to resemble the textual embedding regarding these concepts from the model’s perspective. As the knowledge of the concepts has been transferred to the prompt, this stage helps to reduce the dependency of generating these concepts on current textual inputs. In the knowledge removal stage, assisted by the additional prompt, when eliminating undesirable concepts, the model’s parameters undergo fine-tuning more stably and have a minimal negative impact on other concepts.
4.2 Knowledge Transfer and Removal with Prompt - KPOP
Let us introduce the terminologies in our paper. Notably, we denote $c$ as the input description, commonly referred to as the textual prompt in text-to-image generative models. Additionally, we use $\mathbf{p}$ to represent the parameter prompt, which we will refer to as the ‘prompt’ for brevity henceforth.
Let $\epsilon_{\theta}(z_{t},c,t)$ denote the output of the pre-trained foundation U-Net model parameterized by $\theta$ at step $t$ given an input description $c$ and the latent vector from the previous step $z_{t}$. Let $\epsilon_{\theta^{{}^{\prime}}}(z_{t},c,t)$ denote the output of the sanitized model, parameterized by the to-be-finetuned parameters $\theta^{{}^{\prime}}$. For the sake of simplicity, we employ the notations $\epsilon_{\theta}(c)$ and $\epsilon_{\theta^{{}^{\prime}}}(c)$. Let $\epsilon_{\theta^{{}^{\prime}}}(c,\mathbf{p})$ represent the output of the sanitized model given the input description $c$ and the parameter prompt $\mathbf{p}$. The mechanism that allows us to inject the prompt $\mathbf{p}$ in to the cross-attention layer will be discussed later in Section 4.3. It is worth noting that we can inject the same prompt into multiple cross-attention layers, thus $\epsilon_{\theta^{{}^{\prime}}}()$ denotes the output of the entire model, not just a specific layer. Let $c_{e}\in\mathbf{E}$ denote a textual description in a set of to-be-erased concepts $\mathbf{E}$ and $c_{n}$ represents a neutral or null concept, i.e., ‘a photo’ or ‘ ’.
Knowledge Transfer.
Initially, we initialize $\theta^{\prime}_{0}=\theta$ (i.e., the pre-trained foundation model) and $\mathbf{p_{0}}$. At the iteration $k$, we yield $\theta^{\prime}_{k}$ and $\mathbf{p_{k}}$ and need to update for the next iteration. At this stage, we aim to find $\mathbf{p_{k+1}}$ that is not too far from current $\mathbf{p_{k}}$ and can resemble the undesirable concepts by minimizing the generation loss as (Ho et al., 2020; Song et al., 2020)
$\underset{\mathbf{p}:\|\mathbf{p}-\mathbf{p}_{k}\|_{2}\leq\rho_{p}}{\text{min}}\;\mathbb{E}_{c_{e}\in\mathbf{E}}\left[\left\|\epsilon_{\theta_{k}^{{}^{\prime}}}(c_{e},\mathbf{p})-\epsilon_{\theta}(c_{e})\right\|_{2}^{2}\right].$ (5)
Here, we note that in Eq. 5 as a result of the knowledge removal stage, the model $\epsilon_{\theta^{{}^{\prime}}}$ might have already weakened its knowledge of the undesirable concept, i.e., $\epsilon_{\theta^{{}^{\prime}}}(c_{e})$ differs greatly from $\epsilon_{\theta}(c_{e})$. Therefore, matching $\epsilon_{\theta^{{}^{\prime}}}(c_{e},\mathbf{p})$ with $\epsilon_{\theta}(c_{e})$, i.e., to ensure the fine-tuned model together with the prompt can generate satisfactorily undesirable concepts, allowing the transfer of knowledge of erasing concepts to the prompt. We apply a one-step gradient descent to update the prompt as
$\mathbf{p}_{k+1}=\mathbf{p}_{k}-\eta_{p}\nabla_{\mathbf{p}}\mathcal{L}_{e}\left(\theta^{\prime}_{k},\mathbf{p}\right),$ (6)
where $\mathcal{L}_{e}\left(\theta^{\prime}_{k},\mathbf{p}\right)=\mathbb{E}_{e\in\mathbf{E}}\left[\|\epsilon_{\theta^{\prime}_{k}}\left(c_{e},\mathbf{p}\right)-\epsilon_{\theta}\left(c_{e}\right)\|_{2}^{2}\right]$ and $\eta_{p}$ is the learning rate.
Knowledge Removal.
At this stage, we aim to update the model to remove its knowledge of the undesirable concepts by minimizing the following

### v3

# Hiding and Recovering Knowledge in Text-to-Image Diffusion Models via Learnable Prompts

@@ Abstract
Diffusion models have demonstrated remarkable capability in generating high-quality visual content from textual descriptions. However, since these models are trained on large-scale internet data, they inevitably learn undesirable concepts, such as sensitive content, copyrighted material, and harmful or unethical elements. While previous works focus on permanently removing such concepts, this approach is often impractical, as it can degrade model performance and lead to irreversible loss of information. In this work, we introduce a novel concept-hiding approach that makes unwanted concepts inaccessible to public users while allowing controlled recovery when needed. Instead of erasing knowledge from the model entirely, we incorporate a learnable prompt into the cross-attention module, acting as a secure memory that suppresses the generation of hidden concepts unless a secret key is provided. This enables flexible access control—ensuring that undesirable content cannot be easily generated while preserving the option to reinstate it under restricted conditions. Our method introduces a new paradigm where concept suppression and controlled recovery coexist, which was not feasible in prior works. We validate its effectiveness on the Stable Diffusion model, demonstrating that hiding concepts mitigates the risks of permanent removal while maintaining the model’s overall capability. Our code is available at https://github.com/tuananhbui89/Erasing-KPOP.
@@ 1 Introduction
Recent advances in text-to-image generative models [Rombach et al., 2022, Ramesh et al., 2021, 2022] have garnered significant attention due to their exceptional image quality and creative potential. These models are trained on massive internet-scale datasets, enabling them to generate diverse visual content. However, these datasets often contain undesirable concepts—such as sensitive, harmful, or copyrighted material—that may be inadvertently learned by the models. As a result, users can potentially exploit these models to generate content that spreads misinformation, hate speech, or other unethical material [Rando et al., 2022, Qu et al., 2023, Westerlund, 2019]. To ensure the safety and responsible deployment of these models, it is crucial to prevent the generation of such undesirable content.
To address this challenge, prior works have explored various strategies, including dataset filtering [StabilityAI, 2022], post-generation filtering [Rando et al., 2022], and inference-time guidance [Schramowski et al., 2023]. However, each of these approaches has significant limitations. Specifically, dataset filtering requires extensive human effort to curate training data and has been shown to degrade model performance on general tasks [StabilityAI, 2022], while post-generation filtering (e.g., NSFW classifiers [StabilityAI, 2022]) struggles with false positives and can be easily bypassed by users. Inference-time guidance provides a lightweight solution but can be overridden by adversarial attacks [Gandikota et al., 2023].
A more effective approach is to modify the model itself to remove unwanted concepts. Existing methods [Gandikota et al., 2023, Kumari et al., 2023, Orgad et al., 2023, Zhang et al., 2023, Bui et al., 2024, 2025] attempt to erase knowledge of undesirable concepts through fine-tuning, optimizing specific loss functions to suppress them. However, this permanent removal of concepts has inherent drawbacks. Firstly, cince model parameters are highly shared across concepts, erasing one concept can unintentionally degrade related capabilities. Secondly, once a concept is removed, recovering it requires retraining the model, which may not be feasible in real-world scenarios.
In this work, we introduce a fundamentally different approach: rather than permanently removing, we propose to hide undesirable concepts within the model. Our method prevents public users from generating unwanted content while allowing controlled recovery using a secret key. This paradigm offers key advantages over previous approaches, such as flexibility and controlled access where hidden concepts can be reinstated as needed, enabling adaptive moderation.
Inspired by prompt-based tuning techniques [Li and Liang, 2021, Lester et al., 2021, Pfeiffer et al., 2020], we introduce a learnable prompt to hide undesirable concepts. This prompt serves as an additional memory module that captures knowledge of undesirable concepts, reducing their dependence on core model parameters. High-level speaking, our method consists of two stages: knowledge recovery/transfer - to transfer knowledge of undesirable concepts to the prompt, and knowledge hiding/removal - to hide the undesirable concepts from the model.
Through extensive experiments, we demonstrate that our approach successfully prevents the generation of undesirable content while maintaining the model’s general capabilities. We evaluate this across three scenarios: ➀ Hiding object-oriented concepts, ➁ Mitigating unethical content, and ➂ Hiding artistic style concepts. Our results highlight the practical impact of concept hiding, paving the way for more adaptive and secure content moderation in text-to-image models.
@@ 2 Related Work
In this section, we provide an extensive overview of existing literature concerning concept erasure and related techniques to our work.
@@
Pre-processing methods represent a straightforward approach to eliminating undesired concepts from input images. This involves employing pre-trained detectors to identify images containing objectionable content and subsequently excluding them from the training set. However, the drawback lies in the necessity of retraining the model from scratch, which proves computationally expensive and impractical for evolving erasure requests. A notable instance of complete retraining is evident in Stable Diffusion v2.0 [StabilityAI, 2022], but this approach was reported to leave the model inadequately sanitized [Gandikota et al., 2023].
Post-processing methods encompass the utilization of Not-Safe-For-Work (NSFW) detectors to identify potentially inappropriate content in generated images. Images flagged by the NSFW detector are then either blurred or blacked out before being presented to users. This method, employed by organizations such as OpenAI (developer of Dall-E), StabilityAI (developer of Stable Diffusion), and Midjourney Inc (developer of Midjourney), is considered highly effective. However, the open-source nature of the Stable Diffusion model exposes it to potential evasion by modifying the NSFW detector in the source code. Closed-source models, like Dall-E, are not immune either, as demonstrated in [Yang et al., 2024], where a technique similar to Boundary Attack [Brendel et al., 2017] was used to uncover adversarial prompts that could bypass the filtering mechanism.
Concept Mimicry serves as a personalization technique, generating images aligned with a user’s preferences based on their input. Noteworthy methods include Textual Inversion [Gal et al., 2022] and Dreambooth [Ruiz et al., 2023], which have proven effective with minimal user input. In contrast, Anti Concept Mimicry is employed to safeguard personal or artistic styles from being copied through Concept Mimicry. Achieved by introducing imperceptible adversarial noise to input images, this technique can deceive Concept Mimicry methods under specific conditions. Recent contributions such as Anti-Dreambooth [Van Le et al., 2023] have explored and demonstrated the effectiveness of this approach.
To date, the most successful strategy for sanitizing open-source models, such as Stable Diffusion, involves cleaning the generator (e.g., U-Net) in the diffusion model post-training on raw, unfiltered data and before public release. This approach, as partially demonstrated in Gandikota et al. [2023], underscores the importance of addressing potential biases and undesired content in models before their deployment.
@@
Existing erasing methods [Gandikota et al., 2023, Orgad et al., 2023, Zhang et al., 2023, Kumari et al., 2023] aim to erase undesirable concepts by fine-tuning foundation models with appropriate losses to unlearn and erase these undesirable concepts. Specifically, TIME [Orgad et al., 2023], UCE [Zhang et al., 2023], Concept Ablation [Kumari et al., 2023], and SDD try to project meaning of harmful context into another benign one, while ESD [Gandikota et al., 2023] uses the principle of classifier-free-guidance to remove the distribution of the bad concept from the LDM.
@@
The overarching concept behind prompting involves applying a function to alter the input text, providing the language model with supplementary task-related information. However, devising an effective prompting function presents challenges and necessitates heuristic approaches. Recent studies, such as prompt tuning [Lester et al., 2021] and prefix tuning [Li and Liang, 2021], attempt to tackle this challenge by employing trainable prompts within a continuous space, resulting in impressive performance in transfer learning tasks. Prompts encapsulate task-specific knowledge with significantly fewer additional parameters compared to competing methods like Adapter [Pfeiffer et al., 2020, Wang et al., 2021] and LoRA [Hu et al., 2022].
@@ 3.1 Motivation
Unlearning unwanted concepts, though initially explored, has recently gained significant attention, with several notable works [Gandikota et al., 2023, 2024, Zhang et al., 2023, Bui et al., 2024, 2025] demonstrating that it is possible to remove specific concepts from foundation models before their deployment.
However, in this work, we pursue a different objective compared to existing approaches. Instead of permanently removing unwanted concepts, we aim to hide them—preventing public users from generating undesirable content while allowing model developers to later recover these concepts using a secret key known only to them.
This new hiding paradigm is particularly important in real-world applications for several reasons:
- Flexibility & Controlled Access: Unlike methods that permanently remove concepts, hiding allows model developers to reinstate specific concepts if regulations evolve, new use cases arise, or controlled access is required. This approach also enables the development of novel business models where different levels of access can be granted.
- Enhancing Security Against Backdoor Attacks: Our method can be viewed as a deliberately introduced, controlled backdoor for beneficial purposes. By studying the feasibility of embedding and later recovering hidden concepts, developers can gain insights into potential security vulnerabilities in text-to-image (T2I) models and proactively implement safeguards.
3.2 Knowledge Hiding and Recovery with Prompt - KPOP
High-level Overview
Given a pre-trained foundation model $\theta$ and a set of undesirable concepts $\mathbf{E}$, we propose a fine-tuning process to obtain a sanitized model $\theta^{{}^{\prime}}$ that satisfies the two key objectives: hiding - the undesirable concepts are concealed from $\theta^{{}^{\prime}}$ preventing public users from generating them, and recovery - the hidden concepts can be reinstated when a secret key is provided.
Mathematically, let $\epsilon_{\theta}(z_{t},c,t)$ denote the output of the pre-trained foundation U-Net model parameterized by $\theta$ at step $t$ given an input description $c$ and the latent vector from the previous step $z_{t}$. Similarly, let $\epsilon_{\theta^{{}^{\prime}}}(z_{t},c,t)$ denote the output of the sanitized model, parameterized by parameters $\theta^{{}^{\prime}}$. For simplicity, we employ the notations $\epsilon_{\theta}(c)$ and $\epsilon_{\theta^{{}^{\prime}}}(c)$. The hiding goal can be formulated as $\epsilon_{\theta^{{}^{\prime}}}(c)\neq\epsilon_{\theta}(c)\;\forall c\in\mathbf{E}$, while the recovery goal can be represented as $\epsilon_{\theta^{{}^{\prime}}}(c,\mathbf{p}_{c_{e}})\approx\epsilon_{\theta}(c)\;\forall c\in\mathbf{E}$, where $\mathbf{p}_{c_{e}}$ is a learnable prompt acting as the secret key to recover the concept $c_{e}\in\mathbf{E}$.
To achieve the two objectives, we propose a fine-tuning process that alternates between two stages: knowledge recovery/transfer and knowledge hiding/removal integrated within the following optimization problem: